find the value of n in the following; nP4=30nP2
Answer by ikleyn(52778)

.
The condition gives this equation


n*(n-1)*(n-2)*(n-3) = 30*n*(n-1)  ====>


(n-2)*(n-3) = 30


The product of two consecutive positive integers is  30  ====>  the integers are

n-2 = 6,  n-3 = 5.


Hence n = 8.


Answer.  n = 8.